A camera lens used for taking close-up photographs has a focal length of 25 mm. The farthest it can be placed from the film is 44 mm. What is the magnification of this closest object?

answer to 3 decimal places


1
Expert's answer
2021-06-07T04:51:01-0400

Answer:-

"m=-\frac{d_i}{d_o}\\
d_i=44 \ mm" (distance between the image and the lens )

"d_o=" ? (distance between the object and the lens )

f=25mm ( focal length)

"\frac{1}{d_0}=\frac{1}{f}-\frac{1}{d_i}"

"\frac{1}{d_0}=\frac{1}{25}-\frac{1}{44}"

"d_o=57.894 \ mm"

magnification of this closest object

"m=-\frac{d_i}{d_o}\\"

"m=-\frac{44}{57.894}\\" ="\boxed{0.760}" answer